noobpro78:

Part D The formula for the volume of a cylinder is v=pi r^2 h . What is the approximate volume of two refraction cups? Use 3.14 for pi.

jhonyy9:

what you know about a refraction cup ? dimensions like radius and height these you need to know to can calcule the volume

noobpro78:

idk i suck at math

jhonyy9:

@xxdeppressionxx

noobpro78:

soooooo was that the answer

jhonyy9:

no please make and post a screenshot about your problem

noobpro78:

that in my problaem its based off of the picture

jhonyy9:

so on this image given the length of diameter equal 8 cm from this result the radius equal how many cm ?

xxDeppressionxx:

Since both objects are assumed to have the same volume, matching the formulas we'll get: (πr^2h1/3πr^2h) (r^2=1/3r^2) ( 2√3r^2 = R)

noobpro78:

thx

xxDeppressionxx:

Np